the second angle of a triangle is three times the first and the third is 12 degrees less than twice the first. Find the measures of the angles
Answer by checkley77(12569)

x+3x+(2x-12)=180
6x=180+12
6x=192
x=192/6
x=32 degrees for the first angle.
3*32=96 degrees for the second.
2*32-12=64-12=52 for the third.
32+96+52=180
180=180